OpenStudy (anonymous):

So, i have a triangle and i'm trying to get the length of the side. I need help with this.

OpenStudy (anonymous):

You better type the entire question here.

OpenStudy (saifoo.khan):

Alright.

OpenStudy (saifoo.khan):

post a photo here.

OpenStudy (anonymous):

a^2 + b^2 = c^2 where c is the longest side (hypotenuse) Hope this helps. http://www.tutorsean.net

OpenStudy (anonymous):

What is the length of MN? Triangle LMN with L as a 50 degree angle, N is a 50 degree angle. Side ML shows 4x, and side MN shows x + 3.

jimthompson5910 (jim_thompson5910):

Use the law of sines formula \[\frac{\sin(50)}{x+3}=\frac{\sin(50)}{4x}\] \[\frac{1}{x+3}=\frac{1}{4x}\] \[4x=x+3\] \[4x-x=3\] \[3x=3\] \[x=\frac{3}{3}\] \[x=\frac{3}{3}\] \[x=1\] So the length of MN is \[x+3=1+3=4\] units

OpenStudy (anonymous):

Thank you! :)